the cashier of a fastfood restaurant claims that the average amount spent by customers for dinner is P120. A sample of 50 customers over a month period was randomly selected and it was found out that the average amount spent for dinner was P122.50. Using a 0.05 level of significance, can it be concluded that the average amount spent by customer was more than P120? Assume that the population standard deviation is P6.50. (Use the low-value method)


1
Expert's answer
2021-05-07T12:05:47-0400

Let "H_o" : It can be concluded that the average amount spent by customer is 120


"n=50, \mu=120, x=122.50, \sigma=6.50"


Using z-distribution statistics-


"z=\dfrac{(x-\mu)}{\sigma}=\dfrac{(122.5-120)}{6.5}=0.384"



The tabulated value of z at "\alpha=0.05 1.645"


Conclusion: The calculated value of z is less than the tabulated value so "H_o" is accepted, i.e. It can be concluded that the average amount spent by costumer was 120.